I need help on this problem. Thank you.
Simplify the expression. Must not contain trig function.
tan{(cos^-1)(x/3)}

Answer by lwsshak3(11628)

Simplify the expression. Must not contain trig function.
tan{(cos^-1)(x/3)}
This reads: tan of an angle whose cos=x/3
let A=angle
cosA=1/3
sinA=√(1-cos^2A)=√1-1/9)=√(8/9)=√8/3
tan{(cos^-1)(x/3)}=tanA=sinA/cosA=√8=2√2